Circular No. 6032

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) SL9 message-center reports note that at the Keck Observatory I. de

Pater et al. confirmed the impact of “lost fragment M = 10” (cf. IAUC 6030) around July 20.256 UT, the impact site being clearly separated from that of K, which came into view a little later. Negative reports of K-band observations near the expected time of the P = 8 impact were received from the South African Astronomical Observatory (two groups), the Teide Observatory and Pic du Midi. The Pic du Midi team reported a very short, tiny flash in the K band on July 20.822 that may have been due to fragment Q_{2} = 7b, followed by a double flash from Q_{1} = 7a on 20.842 and 20.847, the latter being the more prominent. Plume phenomena were observed at the Calar Alto Observatory around the impact times of the Q_{2} and Q_{1} fragments; two spot-like features appeared over the impact limb of Jupiter around July 20.822 and 20.847, the latter considerably brighter than the former, and they were simultaneously identified by direct imaging at 2.3 and 1.7 μm and by fast photometry in the 3.1-μm range. A further spot appeared over the limb around July 20.842, and at 2.3 μm it was intermediate in brightness between the others. At the Teide Observatory the Q_{2} plume was noted on July 20.828 (evidently with the IAC-80 telescope at 750, 892 and 948 nm); the impact of Q_{1} was difficult to separate from the Q_{2} plume, but it was clearly observed at 892 nm by July 20.851, the impact cloud later being clearly double. The Q_{2} impact did not show in the raw 2.166-μm data from the Carlos Sanchez telescope, but the Q_{1} flare appeared on July 20.847.

The impact of fragment R = 6 was detected at 2.34 μm using CASPIR on the Australian National University (not Observatory, as was stated on IAUC 6030) 2.3-m telescope at Siding Spring; distinct brightening began on July 21.237 UT and continued until the detector saturated on 21.238; bright emission at the impact site was detected at 4.78 μm on 21.242 but had faded by 21.256. P. Nicholson and others using the 5-m reflector at Palomar first noted the R flare at 3.6 and 5 μm on July 21.233; rapid brightening occurred on 21.237, reaching a peak on 21.240, when diffraction spikes were observed at 5 μm. The 5-μm flux dimmed by a factor of 40 by 21.243, and the image was observed to be extended ∼ 2′′ along the limb on 21.246; by 21.249 the 5-μm flux had dropped to the level of the first images, and by 21.258 it was comparable to that of the planet. At Table Mountain Observatory (Wrightwood, CA) J. Young and others remarked on the appearance (at 890 nm) of the R impact site on Jupiter’s disk on July 21.258, followed by the Q impact site on 21.274.

1994 July 21 Brian G. Marsden